OpenStudy (anonymous):

Please help! (will give a medal) Write a point slope equation for the line that passes through the point (8,1) and is parallel to the line given by y=5.8x-9. Do not use parenthesis on the y side.

OpenStudy (anonymous):

Parallel means that the slope is the same. So the new slope will be 5.8. Now, you can plug in the values you have (8 for x and 1 for y) into y = mx + b. and solve for b. you get: y = mx + b 1 = 5.8(8) + b 1 = 46.4 + b -45.4 = b So your equation is: y = 5.8x-45.4

OpenStudy (anonymous):

y-1=5.8(x-8)

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

I don't get how they get that answer.

OpenStudy (anonymous):

I wrote it in slope-intercept form. They used point slope form: (y-a) = m(x-b) m is the slope of the line (b,a) is a point on the line.

OpenStudy (anonymous):

In the equation y=9x+41 the slope is 9 right?

OpenStudy (anonymous):

That is perfectly correct.

OpenStudy (anonymous):

thanks allot! You're the first person who actually helped me! haha

OpenStudy (anonymous):

glad to be of service.